Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 in the present but Y = 100 in the future. Let MC = 0. Let reserves = 200. Consider the basic two-period model. What is consumption of the resource in the present?

48.32

55.23

62.44

70.12

none of the above

Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 across both periods. Let MC = 30. Let reserves = 200. Consider the basic two-period model. What is consumption of the resource in the present?

92.86

107.14

154.76

200

none of the above
Answer #1

Part 1) We have the following information

Q = 150 – P + 2Y; where Y = 50 in period 1 and Y = 100 in period 2

So, Q1 = 150 – P + 100

Q1 = 250 – P in period 1

Q2 = 150 – P + 200

Q2 = 350 – P in period 2

In the above P is price and Q is the quantity

Marginal cost (MC) = 0

Total reserves = 200

Discount rate (r) = 10% or 0.1

In the case of indefinite stock, the amount of the resources consumed is determined by the profit maximizing condition of Price = MC. However, in the present case the stock is limited to 200. In a two-period model for an efficient allocation we need to equalize the marginal rent of period one and period two.

P – MC(Q1) = P – MC(Q2)

Now, the marginal profit of a given period has to be equal to the discounted marginal profit of the following period. This is called the “r-percent” rule or Hotelling’s rule and it is given by the following equation

P – MC(Q1) = [P – MC(Q2)]/(1 + r)

What the above equation says is that the marginal profit of a given period has to be r% higher than the marginal profit of the previous period. Solving the above

250 – Q1 – 0 = [350 – Q2 – 0]/(1 + 0.1)

250 – Q1 = (350 – Q2)/(1.1)

(250 – Q1) = 350/1.1 – Q2/1.1

Now the reserve constraint is

Q1 + Q2 = 200

We will replace Q2 = 200 – Q1

(250 – Q1) = 350/1.1 – (200 – Q1)/1.1

(250 – Q1) = 318.18 – 200/1.1 + Q1/1.1

(250 – Q1) = 318.18 – 181.81 + Q1/1.1

275 – 1.1Q1 = 150 + Q1

2.1Q1 = 125

Consumption of the Resources in the present: Q1 = 59.52

Q2 = 200 – Q1

Consumption of the Resources in the next period: Q2 = 140.48

So, the correct option is None of the Above.

Part 2) We have the following information

Q = 150 – P + 2Y; where Y = 50 for both the periods

So, Q = 150 – P + 100

Q = 250 – P

P = 250 – Q Where P is price and Q is the quantity

Marginal cost (MC) = 30

Total reserves = 200

Discount rate (r) = 10% or 0.1

In the case of indefinite stock, the amount of the resources consumed is determined by the profit maximizing condition of Price = MC. However, in the present case the stock is limited to 200. In a two-period model for an efficient allocation we need to equalize the marginal rent of period one and period two.

P – MC(Q1) = P – MC(Q2)

Now, the marginal profit of a given period has to be equal to the discounted marginal profit of the following period. This is called the “r-percent” rule or Hotelling’s rule and it is given by the following equation

P – MC(Q1) = [P – MC(Q2)]/(1 + r)

What the above equation says is that the marginal profit of a given period has to be r% higher than the marginal profit of the previous period. Solving the above

250 – Q1 – 30 = [250 – Q2 – 30]/(1 + 0.1)

220 – Q1 = (220 – Q2)/(1.1)

(220 – Q1) = 220/1.1 – Q2/1.1

Now the reserve constraint is

Q1 + Q2 = 200

We will replace Q2 = 200 – Q1

(220 – Q1) = 220/1.1 – (200 – Q1)/1.1

(220 – Q1) = 200 – 200/1.1 + Q1/1.1

(220 – Q1) = 200 – 181.81 + Q1/1.1

242 – 1.1Q1 = 20 + Q1

2.1Q1 = 222

Consumption of the Resources in the present: Q1 = 105.71

Q2 = 200 – Q1

Consumption of the Resources in the next period: Q2 = 94.28

So, the correct option is None of the Above.